Subject: APO 3.5m users committee minutes, August 2011
From: Suzanne Hawley
Submitted: Tue, 6 Sep 2011 15:13:12 -0700 (PDT)
Message number: 1242
(previous: 1241,
next: 1243
up: Index)
APO 3.5-m Users Committee Phonecon, 8/8/11
Attending: Suzanne Hawley, Mark Klaene, Bill Ketzeback, Bob O'Connell
(for Remy Indebetouw), Jon Holtzman, Michael Strauss, Scott Anderson, Sean
Moran, and Bruce Gillespie
Absent: Al Harper, John Bally, Remy Indebetouw, Russet McMillan
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - Scott mentioned the recent question about
drifts in the echelle wavelength stability (see topic discussed below).
o Colorado (John Bally) - No report.
o Johns Hopkins (Sean Moran) - Sean had nothing new to report.
o New Mexico State (Jon Holtzman) - Jon had nothing new to report.
o Princeton (Michael Strauss) - Michael had nothing new to report.
o Virginia (Bob O'Connell) - Bob had nothing new to report.
o Chicago (Al Harper) - No report.
**********************************
Discussion of telescope/instruments report:
Mark Klaene and Bill Ketzeback gave an overview of the past month's
activities. The July summer shutdown is over. The primary mirror was
realuminized, and got a good coating. The echelle upper tank was opened
for maintenance. The DIS dewars were pumped, and power supplies were
replaced. The NA2 instrument rotator was serviced. With the help of
student and temporary staff, we kept schedule; the telescope is up and
running well.
The new camera for the Shack-Hartmann instrument was installed and
commissioned. DIS vacuum and detector temperatures are a little higher
than we like, but DIS seems to be working fine. The blue 400 grating for
DIS acquired a chip during handling. It is not in the beam so the grating
is in service, but we plan to replace it soon. SPIcam, Agile, TripleSpec,
and the echelle are all operational, but the echelle slit viewer camera
computer is crashing. It is an old computer that needs to be upgraded.
The ThAr calibration lamp for the echelle was replaced and the new lamp
has a quartz envelope that attenuates the UV lines -- we will get a
replacement lamp. There have been reports of jumps in the calibration
lines, which is under investigation. The NIC-FPS upside-down problem was
addressed by replacing all the cables in the dewar, and the Leach
controller was replaced to fix excessive read noise. The upgraded GFP is
still in commissioning.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
6/2/11 through 8/2/11
0) Overview
The forest closure was lifted on July 21st after receiving a few inches of
much-needed rain. Some fire restrictions remain in effect. The to-date
rainfall at the observatory is still below normal -- only 4 inches so far.
During the long, dry spring and early summer, numerous fires occurred
around us. All fires within 5 miles of the observatory were kept to under
one acre. The summer shutdown started on July 5th and ended on time, with
July 28th being the first night of science observing.
A) Summer Shutdown
The main task was to re-aluminize the primary mirror at KPNO. Despite a
major detour caused by road construction just outside Kitt Peak, the
mirror arrived at KPNO on July 11 and returned to APO on July 15. Many
thanks to the Starfire folks in Albuquerque for the loan of their shipping
crate again. With the mirror removed, the mirror cell was serviced and
the PMSS re-calibrated. Additional work included: Echelle upper-tank
work and pumping, DIS camera pumping, new DIS robotic power supplies,
top-end metrology, and Na2 rotator bearing service.
The shutdown went quite quickly and smoothly, in part due to additional
temporary manpower. This included 2 UW students (Conor Sayres and Nick
Hunt-Walker) as well as a temporary APO summer helper, Tom Yuhosz. Many
thanks to them.
1) Telescope
The telescope is up and fully operational following the shutdown.
The new Shack-Hartmann camera is now working after considerable effort by
Fritz and Russell to get the new camera into TUI and also by Bill to get
the data-reduction procedure running. It was used to confirm collimation
of the telescope after the mirror was replaced in the mirror cell.
2) Instruments
DIS became quite troublesome during the shutdown. New vacuum gauges that
would have been a great addition to the cameras proved to leak and had to
be removed. Last-minute pumping was needed to get reasonable vacuum
levels. Both Cryotigers had to be purged and recharged. Despite very
good vacuums and nearly perfectly charged cryotiger systems, the cameras
are running 5-10 Deg C warmer than before shutdown. Upon initial testing
the blue-medium grating was found to to be chipped. There are known
deficiencies with the mounting of the gratings in the holder, and some
were corrected but other more invasive fixes were not. The grating has
been glued back together and is serviceable (the chip is outside FOV). A
new grating has been ordered. Other robotics issues were traced to
corroded connections and appear to be OK after cleaning. However, issues
with grating initial calibrations are being investigated (see agenda
item).
SPICam operated with no new issues. Agile is operational; a recent motor
error is being investigated. TripleSpec is operational. Echelle is
operational. Problems with the slitviewer are still appearing. This is
due to an old computer that needs to be upgraded, which is in work but may
take a substantial amount of time. If we cannot get the proper libraries
built and working with MAC OS9 for this older camera, we will have to look
at a new camera similar to what is being considered for the new DIS
slitviewer camera. A replacement of the ThAr lamp in early June was
mentioned in the minutes for the June meeting. Additional issues have
recently surfaced with wavelength calibration (see agenda item below).
The NIC-FPS problem of spontaneous reset of the Leach controller when
rotated upside down has been addressed with a complete replacement of the
in-dewar cabling. This was accomplished at CU along with replacing the
burned-out heater tape. Upon return to APO the read noise was quite high.
This was traced to bad Leach controller cards that were replaced with
spares. The suspect cards will be sent back for repair. The read noise
is now back at 11 DN with the spare cards. Thanks to Stephane Beland for
all his work on NIC-FPS this year.
GFP upgrade work has resumed on the mountain with engineering time this
week. All of the hardware for direct imaging and IFS mode is in place.
Work continues on commissioning characterization and software development.
The problem with thermal stability has been corrected and a fix is in work
for the problem of image drift with rotation. Robotics software has been
completed and is operational.
**********************************
DIS blue grating:
Suzanne emphasized that since the DIS blue low res grating (B400, the one
with the chip) was removed and reinstalled, users should recalibrate
wavelengths for data taken since the shutdown. If there are any problems
with the data or calibrations, let us know.
**********************************
Echelle wavelength calibration:
The same is true for echelle data taken since the shutdown, and we would
particularly like to hear about unexpected wavelength shifts. Suzanne
pointed out that since the new calibration lamp was installed around 6
June, many lines in the blue and UV are gone or very weak. We should have
a new lamp in a few weeks, hopefully in September. UW observer John
Wisniewski has been doing radial-velocity studies with the echelle, and
reports that he see more scatter in the RV residuals (~2 km/sec), and has
also seen two instances of a calibration wavelength jump of ~15 km/sec.
Echelle users are urged to look at their data carefully and report any
problems. Mark mentioned that we are looking in to replacing the echelle
grating, and Jon Holtzman thinks we may have identified one with higher
throughput.
**********************************
TUI support for old versions of operating systems:
Mark suggested that we make MacOS 10.6 the oldest operating system that
will be supported for TUI. Suzanne asked if anyone knows of users using
older versions of MacOS. We do want to maintain platform flexibility, but
it is difficult to keep TUI consistent with several generations of
operating systems.
===> ACTION: We will plan to phase out TUI support for MacOS 10.4 by the
end of the year, and phase out MacOS 10.5 by next summer. Users committee
members should query their users to see if this causes hardships.
**********************************
PISCO instrument update:
Suzanne said that PISCO is a 4-color dichroic imager that is being built
at CfA and was intended to go on Magellan. If APO contributes some money
and manpower, it is possible the instrument could be brought here. There
will be a meeting at APO in August with CfA personnel to discuss details.
The camera has built-in griz filters, but options are being looked at for
other filters. However, it may be more practical to keep either Spicam or
GFP for narrow band imaging. The camera has a 12 arc minute FOV which
will require some reworking of the NA2 port. It is also a large
instrument; one of our existing instruments will need to be moved or
retired to make room for it. Jon asked what the science case is for this
camera. The new capabilities are that it has a much wider field, with
better sensitivity, and obtains four colors simultaneously. Both
Princeton and UW thought these capabilities were very interesting for
future science programs including deep imaging of stellar populations and
transient followup. ReSTAR has also expressed interest in our having such
an instrument.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: User committee members to try to find users who would be
interested in helping develop exposure time calculators.
STATUS: Closed. UW students have been working with Bill
Ketzeback on exposure time calculators for a class project. Anyone who is
interested in helping should contact Bill directly.
===> ACTION: User Committee members to ask their users if MJD in TAI
would be useful information to have written in the FITS headers.
STATUS: Closed. Michael asked if there is any down side in doing
this. Suzanne stated that all instruments should use TAI. Scott said
that as long as the time records are well-labelled, it doesn't matter.
Suzanne thought that the hub could easily write the times as TAI, ignoring
what the ICCs deliver. Bill said that all the ICCs should have standard
time keywords, but making them consistent is not an easy task. Suzanne
said we will look into exactly what the ICCs are writing, especially if
they are only giving UTC times.
[new actions from this meeting]:
===> ACTION: We will plan to phase out TUI support for MacOS 10.4 by
January 2012, and for MacOS 10.5 by July 2012. Users committee members
should query their users to see if this causes hardships.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 12 Sept
at 8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1242 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO